Please help me with this question. I've been trying to solve this question for days but I still couldn't figure this out. We will have to submit this one on September 11. Thanks........
A freight train makes a regular run of 1300 km according to its present schedule. The train dispatchers want to schedule this run for less than one hour, but the average speed will have to be increased by 2 km per hour. What average speed will be required to make the run on the new schedule?

Answer by stanbon(75887)

A freight train makes a regular run of 1300 km according to its present schedule. The train dispatchers want to schedule this run for less than one hour, but the average speed will have to be increased by 2 km per hour. What average speed will be required to make the run on the new schedule?
-------------
Old schedule DATA;
Distance = 1300 km; rate = x kpm ; time = d/r = 1300/x hrs
-----------------
New schedule DATA:
Disance = 1300 km; rate = x+2 kpm ; time = d/r = 1300/(x+2) hrs
-----------------
EQUATION:
New time < 1 hour
1300/(x+2) < 1
x+2 > 1300
x > 1300+2 kph
x > 1303 kph
=================
Comment: That would more than 800 mph. Are you sure you have stated
the problem correctly? It would be more reasonable to REDUCE
the time by one hour rather than to have the new time be LESS
than one hour. Please check the statement of your problem.
=======================
Cheers,
Stan H.